OpenStudy (anonymous):

The event coordinator asks you to determine how many students participated in the track-and-field day. The total number of students in seventh and eighth grade combined is 584; "5" /"8" of them are seventh graders and "3" /"8" of them are eighth graders. If "4" /"5" of the seventh graders participated in track-and-field day and "7" /"8" of the eighth graders participated, about how many total students participated? Describe the process you used to find your answer.

OpenStudy (dan815):

548 5/8 are 7th grade = 5/8*548=#nmber of 7th graders in total 3/8th are 8th graders = 3/8 *548=#of 8th graders in total 4/5 of Total 7th graders = 4/5*5/8*548 7/8 of total 8th graders = 7/8*3/8*548

OpenStudy (anonymous):

thank you so much

OpenStudy (dan815):

ok uw